MOA-II Obsevation Log Monthly Summary 2007/2
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 4750-5041 | 292 | 1384 | 83040 |
LMC | 1132-1307 | 175 | 951 | 266120 |
SMC | 317-339 | 23 | 43 | 12040 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 490 | 2378 | 361200 (100h20m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
113 | 97 | 77 | 61 | 235 | 44 | 59 | 50 | 197 | 39 | 52 | 41 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
39 | 36 | 35 | 36 | 33 | 33 | 32 | 28 | 26 | 21 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
37 | 37 | 35 | 33 | 33 | 40 | 174 | 168 | 164 | 50 | 42 | 42 |
Field13 | Field14 | Field15 | Field16 |
47 | 49 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
23 | 20 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/2/1: DARK:D2062(120sec) and D2063(150sec) are OK. / night time = 24840(sec)
2007/2/2: No observation / night time = 25020(sec)
2007/2/3: D2067(200sec), D2068(100sec) and D2069(60sec) are OK. / night time = 25260(sec)
2007/2/4: D2071(200sec), D2075(60sec) and D2076(300sec) are OK. / night time = 25500(sec)
2007/2/5: D2078(300sec) and D2079(60sec) are OK. / night time = 25800(sec)
2007/2/6: D2081(300sec) and D2082(60sec) are OK. / night time = 26040(sec)
2007/2/7: No observation. / night time = 26280(sec)
2007/2/8: D2083(300sec) is OK. Helium gas was filled in the compressor No 2. / night time = 26520(sec)
2007/2/9: D2084(300sec) and D2085(15sec) are OK. I took flat images(F1578-F1613). / night time = 26760(sec)
2007/2/10: D2086(300sec) and D2087(60sec) are OK. / night time = 27000(sec)
2007/2/11: No observations. / night time = 27240(sec)
2007/2/12: Good Darks D2093 (300s), D2094 (60s) / night time = 27480(sec)
2007/2/13: D2095(300sec) & D2096(60sec) are OK. / night time = 27720(sec)
2007/2/14: Good darks D2097 (300s), D2098 (60s) / night time = 28020(sec)
2007/2/15: D2101(300sec) is OK. / night time = 28260(sec)
2007/2/16: No good darks. Because chip 2, 3 & 4 are noisy. Problem is fixed, but sky is very bright. / night time = 28500(sec)
2007/2/17: D2113(60sec) & D2114(300sec) are OK. / night time = 28740(sec)
2007/2/18: D2116(300sec), D2117(60sec) / night time = 28980(sec)
2007/2/19: D2119(60sec) & D2120(300sec) are OK. / night time = 29220(sec)
2007/2/20: D2121(300sec), D2122(60sec) / night time = 29460(sec)
2007/2/21: D2123(300s), D2124(60s) / night time = 29700(sec)
2007/2/22: D2129(60sec) is OK. Almost all of B4965 are cloudy images. / night time = 29940(sec)
2007/2/23: D2130(300s), D2131(60s) / night time = 30120(sec)
2007/2/24: D2132(60sec) & D3133(300sec) are OK. / night time = 30360(sec)
2007/2/25: Cryocooler is broken. We cannot observe. / night time = 30660(sec)
2007/2/26: No observation. We are cooling the camera. / night time = 30900(sec)
2007/2/27: Good darks D2136(300) / night time = 31140(sec)
2007/2/28: D2137(300s), D2138(60s) / night time = 31320(sec)
TOTAL NIGHT TIME: 786780(sec)
OPERATING RATE: 58.0 %
GB: 17.6 %
LMC: 38.7 %
SMC: 1.7 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month